Put the equation in the form ax2 + bx + c = 0. Use the caret (^) to enter exponents. For example, enter x2 as x^2.
x2 + 6x + 8 = 3

Answer by Simnepi(216)

Let's rearrange your equation.
we subtract 3 from both sides to get rid of the number on the right hand side of the equals sign
giving
x^2+6x+5=0 (quadratic equations are usually made to equal zero!)
Now...
to solve quadratic equations the first thing is to look for factor pairs of the constant term, in this case that is +5.
factors of +5 are: (remember either these have to be both positive or both negative numbers because negative multiplied by negative gives a positive answer)
plus/minus 1 and 5
Sometimes we can have more than 1 factor pair! but since 5 is prime..
Now we look to see if (any of those) factor pairs can be added to make the coefficient of the x term (that's the number in front of the x) which is +6.
Obviously the factor pair we need are plus 1 and 5 because 1 + 5 = +6
Now we get two brackets and put an x in both and one of the factors in each bracket
Like this
(x+1)(x+5)
making this equal zero means that either one of these brackets could equal zero, that is,
either (x+1)=0 or (x+5)=0
If (x+1)=0 then x = -1
OR
If (x+5)=0 then x = -5
These are the two solutions